Trans-Pacific Lumber runs a mill in the Northwest that produces two grades of lumber, A and B, and a by-product, sawdust. The company chooses to allocate the costs on the basis of the physical quantities method.

Last month, it processed 128,000 logs at a total cost of $353,000. The output of the process consisted of 34,300 units of grade A, 51,300 units of grade B, and 15,300 units of sawdust. The sawdust can be sold for $36,280. This is considered to be its net realizable value, which is deducted from the processing costs of the main products.

Required:

What share of the joint costs should be assigned to grade A and grade B? (Do not round intermediate calculations.)

Product Joint costs
Grade A
Grade B
Answer #1

Joint cost to be allocated = $353,000 - $36,280 = $316,720

Total units = 34,300 + 51,300 = 85,600

Joint cost allocation:

Grade A = $316,720 X 34,300 / 85,600 = $126,910

Grade B = $316,720 X 51,300 / 85,600 = $189,810